OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°.

OpenStudy (anonymous):

I ended with r=3(sqroot2) I'm not sure how to progress since the answer choices are

OpenStudy (anonymous):

(3 square root of 2, 315°), (-3 square root of 2, 135°) (3 square root of 2, 225°), (-3 square root of 2, 45°) (3 square root of 2, 45°), (-3 square root of 2, 225°) (3 square root of 2, 135°), (-3 square root of 2, 315°)

OpenStudy (anonymous):

Im stuck idk how to apply the degrees

OpenStudy (imstuck):

refreshing right now...BRB

OpenStudy (anonymous):

okay

OpenStudy (imstuck):

I am getting one of them to be\[(3\sqrt{2},-\frac{ \pi }{ 4 })\]and I think the other would be\[(-3\sqrt{2},-\frac{ \pi }{ 4 })\]That is the best i can do with trying to remember the conversion, ok?

OpenStudy (imstuck):

or maybe it is\[(3\sqrt{2},-\frac{ \pi }{ 4) } \]or\[(3\sqrt{2},-45)\]I think these are the correct answers, to tell you the truth. One is in radians and the other degrees.

OpenStudy (imstuck):

*\[(3\sqrt{2},-\frac{ \pi }{ 4 })\]

OpenStudy (imstuck):

pi over 4 is also 45 degrees.

OpenStudy (imstuck):

Those are the two ways to write (3,-3) in polar form.

OpenStudy (anonymous):

The angle must be between 0 and 360 according to the qyestion

OpenStudy (anonymous):

The other representation is \[ \left (-3 \sqrt{2},\frac{3 \pi }{4}\right) \]

OpenStudy (anonymous):

The answer is \[ \left (3 \sqrt{2},\frac{7 \pi }{4}\right)\\ and \\ \left(-3 \sqrt{2},\frac{3 \pi }{4}\right) \]

OpenStudy (anonymous):

@IMStuck @eliassaab Thank you. My internet got cut off

OpenStudy (anonymous):

NP